1. What is the key step of redox calculations?

2. What is the relationship between normality and molarity? What is the relationship between equivalent and mole? What is the relationship between formula mass and equivalent mass?
Answer #1

Ouestion 1

Key step of redox calculations is the determination of half reactions (Oxidation half and reduction half reaction) and this is done by observing the change in oxidation state. This change of oxidation state will determine the number of electrons which are being transferred from reducing agent to oxidizing agent. Without it, the equation cannot be balanced. Let us take an example:

In this Iron changes its oxidation state from +2 to +3 and chromium changes from +6 to +3. Now to balance:

Let us take another example in which Manganese is silutaneously oxidised and reduced

Question 2

For this let me introduce n factor. It is the ratio of molar mass to equivalent mass.

In an acid, n factor is its basicity. For example for H2SO4 n=2.

In a base, n factor is its acidity. For example for KOH n=1.

In a salt, n factor is total positive or negative charge. For example  Na2SO4 n=2.

n establishes the required relationships asked in the question.

Normality = Molarity X n

equivalent = mole X n

Formula mass = Equivalent mass X n

For example 1.42 g Na2SO4 dissolved in 1L water. For Na2SO4n=2.

Formula mass= 142 g/mol.

Equvalent mass= 71 g/eq.

mole = 1.42/142 = 0.01 mol.

equivalent = 1.42/71 = 0.02 eq.

Molarity= 0.01/1 mol/L = 0.01 mol/L or 0.01 M.

Normality = 0.02/1 eq/L = 0.02 eq/L or 0.02 N.